# Leveraging Machine Learning for High-Dimensional Option Pricing within the Uncertain Volatility Model

### v1

@@ 5.1.3 Geo-Call spread
Table 5: Geo-Call spread in high dimension, for $\rho_{i,j}=0$ and $M<2^{d}$ variable. Values in brackets represent the computational times, measured in seconds.
The results shown in the Tables 4 and 5 are for unrelated underlying assets. To demonstrate that the GTU algorithm also works well with correlated underlyings, we assumed all correlations equal to $0.50$ or $0.75$. The results are shown in Table 6. In this case, we do not have a benchmark for comparison. However, the obtained values demonstrate a high degree of stability, suggesting their accuracy.
$N=16$ $N=32$ $N=64$
@@
Table 6: Geo-Call spread with correlated underlyings. Values in brackets represent the computational times, measured in seconds.
@@ 5.2 Model with uncertainty on correlation
We now test the GTU algorithm under the assumption that the correlation coefficients $\rho_{i,j}$ are also unknown, let them vary between two bounds $\rho_{i,j}^{\min}=-0.5$ and $\rho_{i,j}^{\max}=0.5$. Evaluating an option in this scenario is more complex for two reasons. Firstly, it is necessary to verify that the correlation matrix $\Gamma$, resulting from varying the coefficients $\rho_{i,j}$ during the optimization procedure, remains positive semidefinite. This verification is implemented directly by using MATLAB’s chol function. Secondly, the number of these coefficients increases quadratically with dimension, leading to a rapid increase in computational cost, making it challenging to obtain results in high dimensions.
@@ 5.2.1 Outperformer spread option with variable correlation
The payoff of this option is given in (5.2). We only consider the case $d=2$ since this is the only case considered by Guyon and Labordère [6]. Moreover, in this particular case it is unnecessary to verify that the correlation matrix $\Gamma$ is positive semidefinite, as this holds true for any permissible value of $\rho_{1,2}$. We propose this particular case because it has also been studied by Guyon and Labordère [6] and thus it is possible to test our algorithm against theirs. In this regard, we can observe that our results reported in Table 7 are slightly higher than the reference value in [6]. This slight difference probably results from the fact that our optimization procedure performs better than Guyon and Labordère’s one in [6] and therefore, it is better able to approach the optimal price in UVM. We also note that the price obtained with a variable correlation parameter, specifically $-0.5\leq\rho_{1,2}\leq 0.5$, is significantly higher than the prices obtained considering a fixed correlation of $\rho_{1,2}=0.5$, $\rho_{1,2}=0$ or $\rho_{1,2}=-0.5$, shown in Table 3. Specifically, the higher result for a fixed correlation is achieved for $\rho_{1,2}=-0.5$ and it is equal to $11.39$, while for a variable correlation, we reach $12.77$. To conclude, it is crucial to emphasize the importance of considering a robust model for correlation among the underlying assets, as pricing results can vary significantly. Therefore, variable correlation is a risk factor that practitioners should not overlook.
$N/P$ $125$ $250$ $500$ $1000$ $2000$ $GL$
$32$ $\underset{\left(129\right)}{12.60}$ $\underset{\left(242\right)}{12.72}$ $\underset{\left(577\right)}{12.72}$ $\underset{\left(1754\right)}{12.73}$ $\underset{\left(2302\right)}{12.74}$ $12.67$
@@
Table 7: Outperformer spread option with variable correlation $\rho_{1,2}$. The column labeled “GL” presents the results from [6] for the same option. Values in brackets represent the computational times, measured in seconds.
@@ 5.2.2 Geo-Outperformer option with variable correlation
$\Psi\left(\mathbf{S}_{T}\right)=\left(\sqrt[d-1]{\prod_{k=2}^{d}S_{T}^{k}}-S_{T}^{1}\right)_{+}$ (5.3)
@@
$\rho_{i,j}=\begin{cases}1&\text{if }i=j\\ -0.5&\text{if i$\neq j$ and}\\ 0.5&\text{otherwise.}\end{cases}\min\left(i,j\right)=1$ (5.4)
@@
The results presented in Table 8 are generally consistent with the benchmark, albeit slightly lower. This discrepancy is likely due to the method used to construct the sets $X_{n}$ required for GPR. Specifically, as far as the benchmark is considered, these sets are constructed using optimal correlation in (5.4), whereas for the general problem with variable correlation, they are determined by using the average correlation (4.8), which is sub-optimal. Consequently, more points are needed, compared to the fixed correlation model, to achieve results of equivalent quality.
$N=32$ $N=64$ $N=128$ $BM$
@@
Table 8: Outperformer geo option with variable correlation. Values in brackets represent the computational times, measured in seconds. The column labeled “BM” shows the benchmark computed by using the GTU algorithm with the optimal correlation parameters.
@@ 5.3 Path dependent option: the Call Sharpe option
We conclude our numerical analysis by discussing an option known as Call Sharpe, previously analyzed by Guyon and Labordére in [6]. This option is particularly interesting as it exemplifies a path-dependent option, where the value depends not only on the current price of the underlying asset but also on the previously realized volatility. Specifically, the payoff at maturity, which depends on a single underlying asset, and on its realized volatility which is computed on a monthly base. To this aim, we assume that the maturity time $T$ is a multiple of $1/12$, defining $N_{m}=12T\in\mathbb{N}$ as the number of months comprising the life of the contract. The realized volatility computed using monthly returns is given by
@@
Following Guyon and Labordére [6], at time $t$, the option value depends on $S_{t}$ and on the two path-dependent variables
@@
Moreover, we stress out that, if $t=\tau_{l}$ for a certain $l$ value, than $A_{t}^{2}=S_{t}$, so we can discard this variate when $t$ is a multiple of $1/12$, that is $12t\in\mathbb{N}$.
@@
We stress out that, as far as this path dependent option is considered, we opt for the Monte Carlo method over the quasi-Monte Carlo method to generate the points $\mathbf{x}^{n,p}$ of $X_{n}$ in (5.5), as the dimension of random simulations equals $N_{m}$, to much to make quasi-Monte Carlo be effective with only a few hundreds of samples.
@@
Moreover, as far as this path dependent option is considered, the elements that identify the market state, namely $S,A^{1}$ and $A^{2}$, represent very different quantities from each other. For this reason, as remarked in a similar framework by Goudènege et al. [5], it is beneficial to use a kernel that considers a different length scale for each predictor, such as the ARD Matern 3/2 kernel $k_{\text{M3/2}}^{\text{ARD}}$ which is defines as follows:
$k_{\text{M3/2}}^{\text{ARD}}(\mathbf{x},\mathbf{x}^{\prime})=\sigma_{f}^{2}\left(1+\sqrt{3\sum\left(\frac{\mathbf{x}_{i}-\mathbf{x}^{\prime}_{i}}{\sigma_{l,i}}\right)^{2}}\right)\exp\left(-\frac{\sqrt{3}}{\sigma_{l}}\sqrt{3\sum\left(\frac{\mathbf{x}_{i}-\mathbf{x}^{\prime}_{i}}{\sigma_{l,i}}\right)^{2}}\right),$
@@
In the specific case study considered in our numerical test, we consider $T=1$, so $N_{m}=12$. Results are reported in Table (9). The benchmark value, computed through a PDE approach described in [6], is $58.4$. In comparison, our more reliable price – obtained for $N=384$ and $P=2000$ – yields a value of $57.93$, much closer to the benchmark than the best price by Guyon and Labordère [6], which is worth $55.55$. Additionally, we emphasize that the relative deviation between our price and the benchmark is less than 1%, indicating a very small error. To conclude, we observe that, in this case, a large number of $N$ time steps is necessary to achieve accurate results. This requirement for many time steps necessitates the use of an algorithm with low computational cost per time step. Our tree method, which in this case requires the evaluation of only two future nodes, proves to be an appropriate choice.
$N\backslash P$ $250$ $500$ $1000$ $2000$ $GL$ $BM$
$24$ $\underset{\left(59\right)}{55.97}$ $\underset{\left(109\right)}{55.94}$ $\underset{\left(373\right)}{56.04}$ $\underset{\left(1325\right)}{56.35}$ $55.55$ $58.4$
$48$ $\underset{\left(98\right)}{59.35}$ $\underset{\left(217\right)}{57.26}$ $\underset{\left(751\right)}{57.14}$ $\underset{\left(2030\right)}{57.32}$
$96$ $\underset{\left(199\right)}{57.74}$ $\underset{\left(556\right)}{57.71}$ $\underset{\left(1580\right)}{57.46}$ $\underset{\left(4494\right)}{57.66}$
$192$ $\underset{\left(386\right)}{59.83}$ $\underset{\left(1016\right)}{57.30}$ $\underset{\left(3233\right)}{58.19}$ $\underset{\left(10114\right)}{57.90}$
$384$ $\underset{\left(872\right)}{58.19}$ $\underset{\left(2566\right)}{58.93}$ $\underset{\left(7315\right)}{58.22}$ $\underset{\left(26228\right)}{57.93}$

### v2

@@ 5.1.3 Geo-Call spread
[…]
NNU $N=16$ $N=32$ $N=64$
$d$ $E$ $100$ $200$ $400$ $800$ $100$ $200$ $400$ $800$ $100$ $200$ $400$ $800$ ME BM
$10$ $\underset{\left(47\right)}{9.49}$ $\underset{\left(82\right)}{9.54}$ $\underset{\left(152\right)}{9.55}$ $\underset{\left(309\right)}{9.55}$ $\underset{\left(92\right)}{9.46}$ $\underset{\left(171\right)}{9.54}$ $\underset{\left(329\right)}{9.54}$ $\underset{\left(645\right)}{9.54}$ $\underset{\left(195\right)}{9.45}$ $\underset{\left(358\right)}{9.52}$ $\underset{\left(678\right)}{9.53}$ $\underset{\left(1310\right)}{9.53}$$\pm 0.02$ $9.55$
$20$ $\underset{\left(80\right)}{9.46}$ $\underset{\left(141\right)}{9.51}$ $\underset{\left(263\right)}{9.51}$ $\underset{\left(501\right)}{9.51}$ $\underset{\left(141\right)}{9.49}$ $\underset{\left(255\right)}{9.52}$ $\underset{\left(483\right)}{9.53}$ $\underset{\left(953\right)}{9.53}$ $\underset{\left(258\right)}{9.46}$ $\underset{\left(463\right)}{9.50}$ $\underset{\left(876\right)}{9.51}$ $\underset{\left(1706\right)}{9.51}$$\pm 0.01$ $9.53$
$40$ $\underset{\left(96\right)}{9.49}$ $\underset{\left(172\right)}{9.51}$ $\underset{\left(325\right)}{9.51}$ $\underset{\left(628\right)}{9.51}$ $\underset{\left(172\right)}{9.50}$ $\underset{\left(316\right)}{9.51}$ $\underset{\left(603\right)}{9.51}$ $\underset{\left(1182\right)}{9.51}$ $\underset{\left(346\right)}{9.49}$ $\underset{\left(638\right)}{9.51}$ $\underset{\left(1224\right)}{9.51}$ $\underset{\left(2393\right)}{9.51}$$\pm 0.01$ $9.51$
$80$ $\underset{\left(161\right)}{9.50}$ $\underset{\left(301\right)}{9.51}$ $\underset{\left(579\right)}{9.51}$ $\underset{\left(1138\right)}{9.51}$ $\underset{\left(311\right)}{9.50}$ $\underset{\left(590\right)}{9.50}$ $\underset{\left(1144\right)}{9.50}$ $\underset{\left(2213\right)}{9.50}$ $\underset{\left(523\right)}{9.51}$ $\underset{\left(1010\right)}{9.51}$ $\underset{\left(1910\right)}{9.51}$ $\underset{\left(3696\right)}{9.51}$$\pm 0.01$ $9.51$
Table 7: Geo-Call spread in high dimension, for $\rho_{i,j}=0$ and $M<2^{d}$ variable. The column labeled “BM” shows the benchmark obtained via one-dimensional reduction. The column labeled “ME” gives the margin of error (95% CI half-width) for NNU results. Values in brackets indicate computational times (in seconds).
Tables 6 and 7 presented numerical results in the case of uncorrelated underlying assets. To assess the robustness of the GTU and NNU algorithms under correlated market conditions, we conducted a new set of simulations in which all pairwise correlations were fixed at either $\rho=0.50$ or $\rho=0.75$. The corresponding results are reported in Table 8.
Although no analytical benchmark is available in this setting, both methods display a high degree of numerical stability across all tested configurations. GTU remains accurate and well-behaved up to dimension $d=10$, even under stronger correlation, and shows consistent convergence as the number of paths increases. Similarly, NNU maintains its precision and efficiency, producing prices that evolve smoothly with the number of training epochs and remain robust as both the correlation and the dimension increase.
The good agreement between the two methods, despite the absence of a reference solution, reinforces confidence in the reliability of the computed prices. These findings confirm that both GTU and NNU generalise well to moderately correlated systems, and suggest that their performance remains stable even in the presence of substantial dependence among risk factors.
GTU $N=16$ $N=32$ $N=64$
@@
NNU $N=16$ $N=32$ $N=64$
$d$ $E$ $100$ $200$ $400$ $800$ $100$ $200$ $400$ $800$ $100$ $200$ $400$ $800$ ME
CASE $\rho_{i,j}=0.50$
$2$ $\underset{\left(53\right)}{10.74}$ $\underset{\left(98\right)}{10.79}$ $\underset{\left(189\right)}{10.80}$ $\underset{\left(369\right)}{10.81}$ $\underset{\left(104\right)}{10.79}$ $\underset{\left(194\right)}{10.88}$ $\underset{\left(373\right)}{10.89}$ $\underset{\left(732\right)}{10.90}$ $\underset{\left(196\right)}{10.83}$ $\underset{\left(366\right)}{10.89}$ $\underset{\left(706\right)}{10.90}$ $\underset{\left(1382\right)}{10.91}$$\pm 0.05$
$5$ $\underset{\left(57\right)}{10.62}$ $\underset{\left(105\right)}{10.65}$ $\underset{\left(202\right)}{10.66}$ $\underset{\left(394\right)}{10.66}$ $\underset{\left(130\right)}{10.61}$ $\underset{\left(246\right)}{10.64}$ $\underset{\left(475\right)}{10.64}$ $\underset{\left(961\right)}{10.65}$ $\underset{\left(222\right)}{10.59}$ $\underset{\left(414\right)}{10.64}$ $\underset{\left(799\right)}{10.65}$ $\underset{\left(1568\right)}{10.66}$$\pm 0.03$
$10$ $\underset{\left(61\right)}{10.48}$ $\underset{\left(121\right)}{10.50}$ $\underset{\left(240\right)}{10.51}$ $\underset{\left(477\right)}{10.51}$ $\underset{\left(121\right)}{10.49}$ $\underset{\left(224\right)}{10.52}$ $\underset{\left(430\right)}{10.53}$ $\underset{\left(843\right)}{10.53}$ $\underset{\left(288\right)}{10.53}$ $\underset{\left(545\right)}{10.57}$ $\underset{\left(1027\right)}{10.58}$ $\underset{\left(1903\right)}{10.58}$$\pm 0.02$
CASE $\rho_{i,j}=0.75$
$2$ $\underset{\left(54\right)}{10.89}$ $\underset{\left(99\right)}{10.96}$ $\underset{\left(190\right)}{10.97}$ $\underset{\left(371\right)}{10.97}$ $\underset{\left(104\right)}{10.95}$ $\underset{\left(195\right)}{11.01}$ $\underset{\left(374\right)}{11.03}$ $\underset{\left(733\right)}{11.03}$ $\underset{\left(262\right)}{10.94}$ $\underset{\left(498\right)}{11.03}$ $\underset{\left(971\right)}{11.05}$ $\underset{\left(1902\right)}{11.05}$$\pm 0.05$
$5$ $\underset{\left(59\right)}{10.79}$ $\underset{\left(107\right)}{10.84}$ $\underset{\left(203\right)}{10.85}$ $\underset{\left(394\right)}{10.85}$ $\underset{\left(117\right)}{10.83}$ $\underset{\left(223\right)}{10.87}$ $\underset{\left(435\right)}{10.88}$ $\underset{\left(859\right)}{10.89}$ $\underset{\left(219\right)}{10.90}$ $\underset{\left(410\right)}{10.94}$ $\underset{\left(792\right)}{10.95}$ $\underset{\left(1555\right)}{10.95}$$\pm 0.03$
$10$ $\underset{\left(63\right)}{10.80}$ $\underset{\left(114\right)}{10.83}$ $\underset{\left(217\right)}{10.84}$ $\underset{\left(424\right)}{10.84}$ $\underset{\left(124\right)}{10.86}$ $\underset{\left(231\right)}{10.89}$ $\underset{\left(442\right)}{10.90}$ $\underset{\left(864\right)}{10.91}$ $\underset{\left(247\right)}{10.92}$ $\underset{\left(464\right)}{10.94}$ $\underset{\left(897\right)}{10.95}$ $\underset{\left(1764\right)}{10.96}$$\pm 0.02$
Table 8: Geo-Call spread with correlated underlyings. The column labeled “ME” gives the margin of error (95% CI half-width) for NNU results. Values in brackets indicate computational times (in seconds).
@@ 5.2 Model with uncertainty on correlation
We now test the GTU algorithm under the assumption that the correlation coefficients $\rho_{i,j}$ are also unknown, let them vary between two bounds $\rho_{i,j}^{\min}=-0.5$ and $\rho_{i,j}^{\max}=0.5$. Evaluating an option in this scenario is more complex for two reasons. Firstly, it is necessary to verify that the correlation matrix $\Gamma$, resulting from varying the coefficients $\rho_{i,j}$ during the optimization procedure, remains positive semidefinite. As for GTU, the verification is performed directly using Matlab’s chol function, whereas for NNU, we follow the penalization approach described in Section 4.4. Secondly, the number of these coefficients increases quadratically with dimension, leading to a rapid increase in computational cost, making it challenging to obtain results in high dimensions.
@@ 5.2.1 Outperformer spread option with variable correlation
The payoff of this option is given in (5.2). We only consider the case $d=2$ since this is the only case considered by [Guyon and Henry-Labordère, 2011]. Moreover, in this particular case it is unnecessary to verify that the correlation matrix $\Gamma$ is positive semidefinite, as this holds true for any permissible value of $\rho_{1,2}$. We propose this particular case because it has also been studied by [Guyon and Henry-Labordère, 2011] and thus it is possible to test our algorithm against theirs. In this regard, we can observe that our results reported in Table 7 are slightly higher than the reference value in [Guyon and Henry-Labordère, 2011]. This slight difference probably results from the fact that our optimization procedure performs better than Guyon and Labordère’s one in [Guyon and Henry-Labordère, 2011] and therefore, it is better able to approach the optimal price in UVM. We also note that the price obtained with a variable correlation parameter, specifically $-0.5\leq\rho_{1,2}\leq 0.5$, is significantly higher than the prices obtained considering a fixed correlation of $\rho_{1,2}=0.5$, $\rho_{1,2}=0$ or $\rho_{1,2}=-0.5$, shown in Table 4. Specifically, the higher result for a fixed correlation is achieved for $\rho_{1,2}=-0.5$ and it is equal to $11.39$, while for a variable correlation, we reach $12.77$ with GTU and $12.80\pm 0.05$ with NNU.
We point out that GTU delivers accurate results even with limited computational effort. In particular, for $N=32$ and $P=250$, GTU achieves a price of $12.72$ – very close to the reference value of $12.67$ in under 300 seconds. By comparison, NNU reaches similar levels of accuracy only at higher training costs: the same value is approximated with $N=64$ and $E=400$ demanding a computation time of about $1000$ seconds. This highlights the competitive performance of GTU in low-dimensional settings, especially when rapid estimation is needed.
To conclude, it is crucial to emphasize the importance of considering a robust model for correlation among the underlying assets, as pricing results can vary significantly. Therefore, variable correlation is a risk factor that practitioners should not overlook.
GTU
$N/P$ $125$ $250$ $500$ $1000$ $2000$
$32$ $\underset{\left(129\right)}{12.60}$ $\underset{\left(242\right)}{12.72}$ $\underset{\left(577\right)}{12.72}$ $\underset{\left(1754\right)}{12.73}$ $\underset{\left(2302\right)}{12.74}$
@@
NNU
$N/E$ $100$ $200$ $400$ $800$ $1600$ ME
$32$ $\underset{\left(144\right)}{11.32}$ $\underset{\left(259\right)}{12.49}$ $\underset{\left(491\right)}{12.61}$ $\underset{\left(962\right)}{12.64}$ $\underset{\left(1905\right)}{12.64}$$\pm 0.05$
$64$ $\underset{\left(276\right)}{11.09}$ $\underset{\left(506\right)}{12.48}$ $\underset{\left(968\right)}{12.70}$ $\underset{\left(1890\right)}{12.74}$ $\underset{\left(3721\right)}{12.76}$$\pm 0.05$
$128$ $\underset{\left(634\right)}{11.11}$ $\underset{\left(1176\right)}{12.58}$ $\underset{\left(2301\right)}{12.76}$ $\underset{\left(4591\right)}{12.79}$ $\underset{\left(9180\right)}{12.80}$$\pm 0.05$
Table 9: Outperformer spread option with variable correlation $\rho_{1,2}$. The value computed by [Guyon and Henry-Labordère, 2011] for the same option is $12.67$. The column labeled “ME” gives the margin of error (95% CI half-width) for NNU results. Values in brackets indicate computational times (in seconds).
@@ 5.2.2 Geo-Outperformer option with variable correlation
$\Psi\left(\mathbf{S}_{T}\right)=\left(\sqrt[d-1]{\prod_{k=2}^{d}S_{T}^{k}}-S_{T}^{1}\right)_{+}.$ (5.3)
@@
$\rho_{i,j}=\begin{cases}1&\text{if }i=j,\\ -0.5&\text{if }i\neq j\text{ and }\min(i,j)=1,\\ 0.5&\text{otherwise}\end{cases}$ (5.4)
@@
The results presented in Table 10 are generally consistent with the benchmark, albeit slightly lower for both the two algorithms. In this regard, the less accurate final result in the table is obtained by NNU in the $5$-dimensional case, where the final value of $12.45$ falls below the benchmark of $12.64$, corresponding to a relative error of $-1.5\%$. Nonetheless, this still represents an entirely acceptable level of accuracy. As far as GTU is considered, this discrepancy is likely due to the method used to construct the sets $X_{n}$ required for GPR. Specifically, as far as the benchmark is considered, these sets are constructed using optimal correlation in (5.4), whereas for the general problem with variable correlation, they are determined by using the average correlation (3.18), which is sub-optimal. Consequently, more points are needed, compared to the fixed correlation model, to achieve results of equivalent quality.
As for NNU, the observed underperformance is likely due to the computational complexity of the problem. For instance, in the case $d=5$, the network is required to output $5$ volatilities and $10$ correlations, for a total of $15$ heterogeneous parameters at each time step. Increasing the network’s capacity may help improve accuracy in this setting; however, this comes at the cost of longer computational times and it does not seem to be convenient given the quality of the proposed results.
GTU $N=32$ $N=64$ $N=128$ BM
@@
NNU $N=32$ $N=64$ $N=128$ ME BM
$d$ $E$ $100$ $200$ $400$ $800$ $1600$ $100$ $200$ $400$ $800$ $1600$ $100$ $200$ $400$ $800$ $1600$
$2$ $\underset{\left(163\right)}{13.33}$ $\underset{\left(302\right)}{13.70}$ $\underset{\left(581\right)}{13.74}$ $\underset{\left(1139\right)}{13.75}$ $\underset{\left(2255\right)}{13.75}$ $\underset{\left(279\right)}{13.19}$ $\underset{\left(513\right)}{13.73}$ $\underset{\left(981\right)}{13.78}$ $\underset{\left(1916\right)}{13.78}$ $\underset{\left(3788\right)}{13.79}$ $\underset{\left(592\right)}{13.10}$ $\underset{\left(1087\right)}{13.65}$ $\underset{\left(2077\right)}{13.69}$ $\underset{\left(4061\right)}{13.70}$ $\underset{\left(7832\right)}{13.70}$$\pm 0.13$ $13.75$
$3$ $\underset{\left(271\right)}{12.32}$ $\underset{\left(511\right)}{12.78}$ $\underset{\left(992\right)}{12.83}$ $\underset{\left(1947\right)}{12.83}$ $\underset{\left(3805\right)}{12.84}$ $\underset{\left(626\right)}{12.16}$ $\underset{\left(1189\right)}{12.77}$ $\underset{\left(2306\right)}{12.83}$ $\underset{\left(4535\right)}{12.84}$ $\underset{\left(8982\right)}{12.84}$ $\underset{\left(1064\right)}{12.16}$ $\underset{\left(2016\right)}{12.77}$ $\underset{\left(3901\right)}{12.81}$ $\underset{\left(7649\right)}{12.83}$ $\underset{\left(15116\right)}{12.84}$$\pm 0.12$ $12.96$
$4$ $\underset{\left(327\right)}{11.57}$ $\underset{\left(620\right)}{12.59}$ $\underset{\left(1197\right)}{12.67}$ $\underset{\left(2344\right)}{12.68}$ $\underset{\left(4636\right)}{12.68}$ $\underset{\left(647\right)}{11.87}$ $\underset{\left(1230\right)}{12.50}$ $\underset{\left(2381\right)}{12.55}$ $\underset{\left(4669\right)}{12.56}$ $\underset{\left(9234\right)}{12.56}$ $\underset{\left(1280\right)}{11.68}$ $\underset{\left(2444\right)}{12.56}$ $\underset{\left(4743\right)}{12.61}$ $\underset{\left(9350\right)}{12.62}$ $\underset{\left(18552\right)}{12.62}$$\pm 0.11$ $12.73$
$5$ $\underset{\left(469\right)}{11.59}$ $\underset{\left(898\right)}{12.33}$ $\underset{\left(1746\right)}{12.37}$ $\underset{\left(3435\right)}{12.38}$ $\underset{\left(6798\right)}{12.38}$ $\underset{\left(1057\right)}{11.68}$ $\underset{\left(2040\right)}{12.40}$ $\underset{\left(3980\right)}{12.45}$ $\underset{\left(7826\right)}{12.46}$ $\underset{\left(15486\right)}{12.46}$ $\underset{\left(1839\right)}{11.83}$ $\underset{\left(3539\right)}{12.40}$ $\underset{\left(6920\right)}{12.44}$ $\underset{\left(13656\right)}{12.45}$ $\underset{\left(26901\right)}{12.45}$$\pm 0.11$ $12.64$
Table 10: Outperformer geo option with variable correlation. Values in brackets represent the computational times, measured in seconds. The column labeled “BM” shows the benchmark computed by using the GTU algorithm with the optimal correlation parameters. The column labeled “ME” gives the margin of error (95% CI half-width) for NNU results. Values in brackets indicate computational times (in seconds).
@@ 5.3 Path dependent option: the Call Sharpe option
We conclude our numerical analysis by discussing an option known as Call Sharpe, previously analyzed by [Guyon and Henry-Labordère, 2011]. This option is particularly interesting as it exemplifies a path-dependent option, where the value depends not only on the current price of the underlying asset but also on the previously realized volatility. Specifically, the payoff at maturity, which depends on a single underlying asset, and on its realized volatility which is computed on a monthly base. To this aim, we assume that the maturity time $T$ is a multiple of $1/12$, defining $N_{m}=12T\in\mathbb{N}$ as the number of months comprising the life of the contract. The realized volatility computed using monthly returns is given by
@@
Following [Guyon and Henry-Labordère, 2011], at time $t$, the option value depends on $S_{t}$ and on the two path-dependent variables
@@
At the monthly monitoring dates $t=\tau_{l}$, it holds by construction that $A^{2}_{t}=S_{t}$. As a result, the variable $A^{2}_{t}$ provides no additional information beyond what is already captured by $S_{t}$, and can thus be considered redundant. Therefore, the pricing problem at such times can be formulated as a function of only two variables – $S_{t}$ and $A^{1}_{t}$ – instead of three. This dimensionality reduction simplifies the regression task and improves computational efficiency.
@@
We emphasize that, as far as this path dependent option is considered, we opt for the Monte Carlo method over the quasi-Monte Carlo method to generate the points $\mathbf{x}^{n,p}$ of $X_{n}$ in (5.5), as the dimension of random simulations equals $N_{m}$, to much to make quasi-Monte Carlo be effective with only a few hundreds of samples.
@@
Moreover, as far as this path dependent option is considered, the elements that identify the market state, namely $S,A^{1}$ and $A^{2}$, represent very different quantities from each other. For this reason, as remarked in a similar framework by [Goudenège et al., 2020], it is beneficial to use a kernel that considers a different length scale for each predictor, such as the ARD Matérn 3/2 kernel $k_{\text{M3/2}}^{\text{ARD}}$ which is defines as follows:
$k_{\text{M3/2}}^{\text{ARD}}(\mathbf{x},\mathbf{x}^{\prime})=\sigma_{f}^{2}\left(1+\sqrt{3\sum_{i}\left(\frac{\mathbf{x}_{i}-\mathbf{x}^{\prime}_{i}}{\sigma_{l,i}}\right)^{2}}\right)\exp\left(-\frac{\sqrt{3}}{\sigma_{l}}\sqrt{3\sum_{i}\left(\frac{\mathbf{x}_{i}-\mathbf{x}^{\prime}_{i}}{\sigma_{l,i}}\right)^{2}}\right),$
@@
Adaptations similar to those discussed for GTU are also necessary for NNU. In addition, given the computational complexity of the problem under consideration, we increased the number of neurons for each layer from $32$ to $64$. Moreover, although the number of independent inputs is reduced to three at monitoring times, we choose to always feed the network with $5$ inputs, namely time $t_{n}$, intra-interval position $\sup_{\{l\mid\tau_{l}\leq t\}}\tau_{l}$ (i.e., the time elapsed since the last monitoring date), current price $S_{n}$, accumulated variance $A^{1}_{t_{n}}$, and the last monitored price $A^{2}_{t_{n}}$. This design ensures that the network architecture remains uniform across all time steps.
In the specific case study considered in our numerical test, we consider $T=1$, so $N_{m}=12$. Results are reported in Table (11). The benchmark value, computed through a PDE approach described in [Guyon and Henry-Labordère, 2011], is $58.40$.
For GTU, the most accurate result is obtained with $N=384$ and $P=2000$, yielding a price of 57.93, which is much closer to the benchmark than the best value reported in [Guyon and Henry-Labordère, 2011] (55.55). The relative deviation is below 1%, indicating a high level of accuracy. This test case also confirms that a large number of time steps $N$ is needed to attain accurate results, thus highlighting the importance of using a low-cost algorithm per time step. The GTU method, which in this case evaluates only two future values per point, proves particularly efficient in this regard.
As for NNU, the results show consistent improvement as both the number of time steps $N$ and the number of training epochs $E$ increase. Although the final estimates remain slightly below the PDE benchmark, the accuracy improves steadily with training, reaching $57.08\pm 0.44$ for $N=384$ and $E=3200$, with a relative deviation of approximately 2.3%. While NNU requires longer computational times than GTU to reach this level of precision, it remains competitive and exhibits stable convergence.
In conclusion, GTU achieves more accurate results at lower computational cost for this problem. However, NNU remains a valid alternative, particularly when flexibility and model generalisation are prioritized, and its performance continues to improve with deeper training.
Remark 9.
The convergence of both GTU and NNU is primarily influenced by two factors: the number of time steps $N$ used in the discretisation of the control problem, and the size of the statistical model (either the number of training points $P$ for kriging in GTU, or the number of epochs $E$ in the training of the neural net in NNU).
In the case of GTU, increasing $N$ improves the resolution of the dynamic programming recursion and reduces the discretisation bias. For path-independent options, numerical tests show that results become stable already at $N\approx 64$ (for instance, in the Geo-Call spread, the estimated price varies by less than $0.04$ between $N=16$ and $N=64$; see Table 6). For path-dependent payoffs, such as the Call-Sharpe contract, a finer time grid is needed, typically $N\geq 256$, to obtain results that match the PDE benchmark within a fraction of a percent (see Table 11). Similar behaviour is observed with NNU, where path-independent contracts require a coarse grid only, while path-dependent options benefit from a finer temporal resolution.
The second key component is the statistical approximation. In GTU, the Gaussian process surrogate is based on a Matérn $3/2$ kernel, whose mean-square error decreases with the number of training points $P$, depending on the input dimension and the smoothness of the target. Empirically, we observe that $P=500$ to $1000$ suffices to make the regression error negligible compared to the discretisation bias. In the NNU, the accuracy depends instead on the number of training epochs $E$. Due to the expressive power of feedforward neural nets [Hornik, 1991] and their effectiveness in hedging applications [Buehler et al., 2019], the pricing error decreases rapidly with $E$. In the Geo-Call example, the NNU estimate lies within the Monte Carlo confidence band after $E\approx 400$, and stabilises around the benchmark value by $E=800$.
GTU NNU
$N/P$ $250$ $500$ $1000$ $2000$ $N/E$ $400$ $800$ $1600$ $3200$ ME
$24$ $\underset{\left(59\right)}{55.97}$ $\underset{\left(109\right)}{55.94}$ $\underset{\left(373\right)}{56.04}$ $\underset{\left(1325\right)}{56.35}$ $24$ $\underset{\left(299\right)}{51.77}$ $\underset{\left(590\right)}{52.64}$ $\underset{\left(1171\right)}{52.91}$ $\underset{\left(2333\right)}{52.97}$$\pm 0.42$
$48$ $\underset{\left(98\right)}{59.35}$ $\underset{\left(217\right)}{57.26}$ $\underset{\left(751\right)}{57.14}$ $\underset{\left(2030\right)}{57.32}$ $48$ $\underset{\left(707\right)}{53.36}$ $\underset{\left(1405\right)}{54.32}$ $\underset{\left(2802\right)}{54.58}$ $\underset{\left(5592\right)}{54.62}$$\pm 0.42$
$96$ $\underset{\left(199\right)}{57.74}$ $\underset{\left(556\right)}{57.71}$ $\underset{\left(1580\right)}{57.46}$ $\underset{\left(4494\right)}{57.66}$ $96$ $\underset{\left(1277\right)}{53.71}$ $\underset{\left(1277\right)}{53.71}$ $\underset{\left(5060\right)}{55.54}$ $\underset{\left(10103\right)}{55.60}$$\pm 0.42$
$192$ $\underset{\left(386\right)}{59.83}$ $\underset{\left(1016\right)}{57.30}$ $\underset{\left(3233\right)}{58.19}$ $\underset{\left(10114\right)}{57.90}$ $192$ $\underset{\left(2821\right)}{55.14}$ $\underset{\left(5628\right)}{55.98}$ $\underset{\left(11240\right)}{56.27}$ $\underset{\left(21109\right)}{56.32}$$\pm 0.43$
$384$ $\underset{\left(872\right)}{58.19}$ $\underset{\left(2566\right)}{58.93}$ $\underset{\left(7315\right)}{58.22}$ $\underset{\left(26228\right)}{57.93}$ $384$ $\underset{\left(6722\right)}{55.18}$ $\underset{\left(12768\right)}{56.41}$ $\underset{\left(12520\right)}{56.66}$ $\underset{\left(24569\right)}{57.08}$$\pm 0.44$